OpenStudy (anonymous):

How to simplify 5 sqrt {3{162}

OpenStudy (anonymous):

\[\Huge 5\sqrt[3]{162} \] ?? :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

162 can you make into a prime numbers? like: 162=81*2 then 81=3*27 so you'd have: 162=3*27*2 and you can write 27 as 3^3 so what would we do with it?? \[\LARGE 5\sqrt[3]{3\cdot 3^3\cdot 2}\] ?

OpenStudy (anonymous):

the whole problem needs to be simplified into something like this 3{3{6}

OpenStudy (anonymous):

\[\LARGE 5\sqrt[3]{3\cdot 3^3\cdot 2}=5\underbrace{\sqrt[3]{3^3}}_{This }\cdot \sqrt[3]{6}\] I want to know what can we do with it..? :)

OpenStudy (anonymous):

15 {3{6}

OpenStudy (anonymous):

yes... well done \[\huge 15\sqrt[3]{6}\]